# An Exploratory Study on Using Large Language Models for Mutation Testing

### v1

@@ 4.1.3. AST Node Dirversity
To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by Javalang (jav, 2024), to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 shows the results, which reveal that LLM-based mutation approaches introduce more types of new AST nodes than other approaches. GPT-3.5 exhibits the greatest diversity, introducing 45 new AST node types, and subsuming all other approaches. Traditional approach Major introduces only two new types of AST nodes, which is not surprising as the mutation operators are a fixed set of simple operations.
We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Few-shot Example” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node. This indicates that the LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
@@
Top K GPT-3.5 CodeLlama-13b LEAM $\mu$BERT Major
Del. 5.1% 0.3% 49.7% 0.2% 17.6%
1st LT (47.6%) LT (49.5%) LT (23.6%) LT (53.8%) ST (59.9%)
2nd BO (24.7%) BO (35.5%) MI (17.6%) MR (22.7%) LT (40.1%)
3rd MI (8.2%) MI (4.1%) MR (16.3%) MI (12.5%)      —
ST: Statement, BO: BinaryOperation, LT: Literal, MI: MethodInvocation, MR: MemberReference
We further analyzed the distribution of deletions and AST Node types across different approaches, as shown in Table 6, which highlights the top-3 most frequent AST node types. For the proportion of deletions, GPT-3.5 and CodeLlama-13b generate 5.1% and 0.3% deletions, respectively. Nearly half of the mutations of LEAM are deletions, which may be the reason for generating many non-compilable mutations. Major generates deletions on restricted locations and results in 17.6% deletions. $\mu$Bert replaces the masked code elements with BERT, thereby rarely having deletions.
Regarding AST Node types, the most frequent new AST node type of all the learning-based approaches is Literal (e.g., $a>b\mapsto\text{true}$). However, for the rule-based approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the Deletion operator, there exists a significant portion of producing empty statements, leading to a large proportion.
Finding 3: LLM-based mutation approaches introduce more types of new AST nodes than other approaches. The mutations generated by GPT-3.5 introduce 45 new types of AST nodes, demonstrating the highest level of diversity.
@@ 4.2.1. Real Bug Detectability
GPT-3.5 detects 382 bugs of the 395 Defects4J bugs and 39 bugs of all 45 ConDefects bugs, i.e., 96.7% Defects4J bugs and 86.7% ConDefects bugs could be revealed by these mutations, achieving the best performance. CodeLlama-13b detects 358 bugs from Defects4J (i.e., 90.6%) and 30 bugs from ConDefects (i.e., 66.7%), respectively. Major achieves the second-best performance by detecting 362 Defects4J bugs (i.e., accounting for 91.6%) and 31 ConDefects bugs (i.e., accounting for 68.9%).
Finding 4: GPT-3.5 has the best performance in Real Bug Detectability, identifying 96.7% of bugs in Defects4J and 86.7% in ConDefects. Among the other approaches, Major performs best, detecting 91.6% of bugs in Defects4J and 68.9% in ConDefects.
@@ 4.2.2. Coupling Rate
The Coupling Rate measures the degree of coupling between the generated mutations and their corresponding real bugs. GPT-3.5 exhibits coupling rates of 0.416 on Defects4J and 0.625 on ConDefects, respectively, achieving the best performance on both datasets. While CodeLlama-13b achieves coupling rates of 0.398 and 0.612, respectively.
Finding 5: GPT-3.5 leads in Coupling Rate with scores of 0.416 on Defects4J and 0.625 on ConDefects. $\mu$Bert follows closely with 0.409 on Defects4J and 0.581 on ConDefects.
@@ 4.2.3. Ochiai Coefficient
The Ochiai Coefficient measures the semantic similarity between mutations and real bugs. GPT-3.5 leads with coefficients of 0.638 on Defects4J and 0.689 on ConDefects, outperforming CodeLlama-13b, which scores 0.39 and 0.378 on the respective datasets. Despite the notable performance gap between the two, their results are consistent across datasets. Major ranks second with coefficients of 0.519 on Defects4J and 0.6 on ConDefects.
Finding 6: GPT-3.5 leads in Ochiai Coefficient performance with scores of 0.519 on Defects4J and 0.6 on ConDefects. Major closely follows, matching these scores and securing the second-best ranking among the approaches.
4.3. RQ3: Impacts of Different Prompts
Table 7. Comparison Results of Different Prompts and LLMs
Metric Category Metric GPT-3.5 Closed Source LLM Open Source LLM
P1 P2 P3 P4 GPT-3.5 GPT-4 StarChat-16b CodeLlama-13b
Mutation Count 47,666 47,905 41,548 39,638 47,666 23,526 25,496 41,808
Mutation Score 0.705 0.703 0.731 0.782 0.705 0.702 0.538 0.734
Cost \$ Cost per 1K Mutations 0.288 0.288 0.287 0.289 0.288 18.701 0.876 0.577
Average Generation Time 1.79 1.78 1.81 1.88 1.79 42.25 13.76 9.06
Usability Compilability Rate 65.7% 64.1% 61.9% 60.8% 65.7% 79.9% 58.2% 74.1%
Useless Mutation Ratio 13.7% 11.3% 22.3% 15.0% 13.7% 7.2% 38.3% 39.2%
Equivalent Mutation Rate 1.5% 1.2% 1.2% 1.5% 1.5% 0.9% 2.1% 1.0%
Behavior Real Bug Detectability 0.895 0.857 0.714 0.629 0.895 0.905 0.324 0.724
Coupling Rate 0.140 0.136 0.096 0.096 0.140 0.143 0.101 0.110
Ochiai Coefficient 0.590 0.571 0.438 0.362 0.590 0.619 0.152 0.286
@@ 4.3. RQ3: Impacts of Different Prompts
The left half of Table 7 presents the comparative results of GPT-3.5 via different prompts listed in Section 3.5.3. Prompts P1 to P3 progressively simplify, each containing less information than its predecessor, while P4 is the most complex, enhancing P1 with test suite codes.
@@
Overall, P1, the default prompt, excels in Compilability Rate and all Behavior Metrics. P2, created by removing few-shot examples from P1, leads in Average Generation Time, Useless Mutation Ratio, and Equivalent Mutation Rate, suggesting improved quality in compilable mutations. P3, provided only with the code element to be mutated, achieves the lowest cost for using the least tokens. Conversely, P4, which extends P1 with test suite codes, shows the lowest performance across all metrics, suggesting that GPT-3.5 cannot effectively utilize test suite data to enhance mutation quality.
@@
Finding 7: The prompt with the whole method and few-shot examples as context (i.e., P1) archives the best performance across all behavior metrics, whereas adding the code of test suites (i.e., P4) decreases the performance.
4.4. RQ4: Impacts of Different LLMs
To answer this RQ, we add two extra LLMs, GPT-4 and StarChat-16b, and compare their results with the two default LLMs, GPT-3.5 and Code Llama-13b. The right half of Table 7 shows comparative results across models using the default prompt. We observe that closed-source LLMs generally outperform others in most metrics. GPT-3.5 excels in Mutation Count, Generation Cost per 1K Mutations, and Average Generation Time, ideal for rapidly generating numerous mutations. GPT-4 leads in all Usability Metrics, and Behavior Similarity Metrics, demonstrating its effectiveness in code-related tasks, although its enhancements over GPT-3.5 in Behavior Metrics are trivial. Between the two open-source LLMs, despite StarChat-$\beta$-16b having more parameters, CodeLlama-13b outperforms in all metrics. This suggests that model architecture and training data quality significantly impact performance beyond just the number of parameters.
Finding 8: Among the four LLMs we study, GPT-3.5 excels in generating a large number of mutations with lower costs, while GPT-4 outperforms in all the Usability Metrics, and Behavior Similarity Metrics.
@@ 4.5. RQ5: Root Causes and Error Types of Non-Compilable Mutations
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 4.1, LLMs generate a significant number of non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT-3.5, ensuring the confidence level is 95% and the margin of error is 5%. From the manual analysis of these non-compilable mutations, we identified 9 distinct error types, as shown in Table 8.
@@
Shown as Table 8, the most common error type, Usage of Unknown Methods, accounts for 27.34% of the total errors, revealing the hallucination issue of generative models (Huang et al., 2023). Code Structural Destruction is the second most common error, accounting for 22.92%, indicating that ensuring generated codes are syntactically correct remains a challenge for current LLMs. This result suggests that there is still significant room for improvement in current LLMs.
Finding 9: The mutations generated by GPT-3.5 have 9 main compilation error types, with Usage of Unknown Methods and Code Structural Destruction being the top-2 most prevalent types.
@@
(b) CodeLlama-13b
(c) LEAM
(d) $\mu$Bert
(e) Major
@@
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by GPT-3.5, CodeLlama, LEAM, and $\mu$Bert in Section 4.1, as shown in Figure 3. For all approaches, the code locations with MethodInvocation and MemberReference are the top-2 prevent AST Node types. In particular, there are more than 30% of non-compilable mutations occur at the location with MethodInvocation, and 20% occur at the location with MemberReference. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that for GPT-3.5, CodeLlama, LEAM, $\mu$Bert, and Major, these mutations account for 7.1%, 0.2%, 45.3%, 0.14%, and 14.4% of all their non-compilable mutations, respectively. Thus for LLMs, deletion is not the major reason for non-compilation.
Finding 10: Method invocations and member references in the code are the most likely to trigger LLMs to generate non-compilable mutations.
@@ 5.1.1. Experiment Setup of the Context Length
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 9. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in Table 10. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the $p$-values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
@@ 5.1.2. Impact of Different Few-Shot Examples
Table 12. Performance Similarity of Different Few-shot Examples
Settings Spea. Coef. $p$-value Pear. Coef $p$-value
QB-6a (Default) v.s. QB-3 0.9701 6.55E-05 0.9996 1.70E-10
QB-6a (Default) v.s. QB-3b 0.9762 3.31E-05 0.9984 1.00E-08
QB-6a (Default) v.s QB-9 0.9762 3.31E-05 0.9985 7.98E-09
QB-6a (Default) v.s D4J-6 0.9524 2.60E-04 0.9980 2.13E-08
@@
We further compute their similarity, shown as Table 12. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the $p$-values are all below 0.05. Therefore, we can conclude the performance of different prompts is very similar.

### v2

@@ 4.1.3. AST Node Dirversity
To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 shows the number of introduced types, which reveals that LLM-based mutation approaches introduce more types of new AST nodes than other approaches. The best of LLMs, GPT-4o-Mini and DeepSeek, exhibit the greatest diversity, introducing 57 new AST node types. Among the traditional approaches, LEAM, $\mu$Bert, and Major introduce 31, 30, and 2 new types of AST nodes, respectively. Additionally, all six LLMs significantly are higher than traditional approaches in AST node types. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
@@
Top K GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert Major
Del. 5.7% 2.2% 0.5% 1.1% 0.2% 0.1% 48.4% 0.6% 20.2%
1st LT(36.1%) LT(43.5%) LT(41.7%) LT(40.1%) LT(20.1%) LT(42.5%) LT(21.4%) LT(25.4%) ST(58.6%)
2nd BO(17.4%) BO(16.5%) BO(22.0%) BO(18.7%) MR(13.7%) BO(23.8%) MR(14.7%) SE(21.7%) LT(41.4%)
3rd MI(10.9%) MI(7.9%) MI(7.1%) MI(10.5%) BO(11.8%) MI(8.0%) MI(12.0%) MR(21.9%) ——
LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 6, which highlights the top three most frequent AST node types. For the proportion of deletions, LEAM shows the highest deletion mutation ratio, with 48.4%, followed by Major at 20.2%. In contrast, all six LLMs generate deletion mutations at a rate below 6%, with three models producing fewer than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we found that deletions are one of the key factors contributing to the lower compilation success rate of traditional approaches.
Regarding AST node types, the most frequent new AST node type of all the learning-based approaches is Literal (e.g., $a>b\mapsto\text{true}$). However, for the rule-based approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
Finding 3 (RQ1): LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches, while less inclined to generation deletion mutations.
@@ 4.2.1. Real Bug Detectability
Among the six LLMs, five models (excluding StarChat) outperform all the traditional approaches. Of the traditional approaches, the rule-based approach Major performs the best, detecting 74.4% of the bugs. GPT-4o and GPT-4o-Mini achieve the highest detection rates, identifying 93.4% of the 851 bugs, excelling Major by 19% on all bugs.
In particular, on the ConDefects dataset, which mitigates data leakage risks, the real bug detection rates for the various approaches are as follows: GPT-3.5 (78.9%), GPT-4o (84.6%), GPT-4o Mini (82.5%), DeepSeek (83.7%), StarChat (32.9%), CodeLlama (63.4%), LEAM (48.9%), $\mu$Bert (47.2%), PIT (49.2%), and Major (56.1%). The best of LLMs (i.e., GPT-4o) outperforms nearly 29% over the best of traditional approaches (i.e., Major), highlighting the significant advantage that LLMs hold over traditional techniques in detecting real bugs.
Finding 4 (RQ2): In Real Bug Detectability, five out of six LLMs significantly outperform traditional approaches. In particular, on ConDefects, GPT-4o outperforms nearly 29% over the best of the traditional approach, Major, highlighting the advantages of LLMs in detecting real bugs.
@@ 4.2.2. Coupling Rate
The Coupling Rate measures the degree of coupling between the generated mutations and their corresponding real bugs.
Among traditional approaches, $\mu$Bert is the best one with a rate of 41.6%. However, two LLMs outperform all traditional approaches in this metric, achieved by the GPT-4 LLMs. The best LLM, GPT-4o-Mini, exhibits a rate of 44.1%, exceeding $\mu$Bert by 2.5%.
Finding 5 (RQ2): GPT-4 achieves the highest Coupling Rate at 44.1%, outperforming all traditional approaches, including the best traditional approach, $\mu$Bert, by 2.5%.
@@ 4.2.3. Ochiai Coefficient
The Ochiai Coefficient measures the semantic similarity between mutations and real bugs (Ojdanic et al., 2023a).
Similar to real bug detectability, most LLMs significantly outperform all existing approaches. Among the LLMs, GPT-4o achieves the highest Ochiai coefficient at 68.5%, 24.1% higher than the best rule-based approach, Major (44.4%), and 31.1% higher than the best small model-based approach, LEAM (37.4%). Moreover, five out of the six LLMs (excluding StarChat), outperforms all traditional approaches.
Finding 6 (RQ2): In the Ochiai Coefficient, five LLMs surpass all traditional generation approaches. In particular, the best LLM, GPT-4o, exceeds the best of traditional approaches, Major, by 24.1%.
4.3. RQ3: Impacts of Different LLMs
As shown in Table 4, we compare the six LLMs we studied in mutation generation. The results reveal significant performance differences among the models.
For Cost, GPT-4o-Mini is the fastest, requiring only 39% of the time taken by the fastest open-source model, DeepSeek.
For Usability, GPT-4o and GPT-4o-Mini lead in all metrics with similar performance, while DeepSeek closely follows and performs similarly to GPT-3.5. Additionally, as shown in Figure 2, GPT-4o-Mini and DeepSeek exhibit the most diverse AST node types, indicating their mutations are more expressive.
For Behavior, GPT-4o and GPT-4o-Mini excel in all metrics, and DeepSeek has similar performance to GPT-3.5, making it the best open-source LLM. StarChat and CodeLlama underperform across most metrics, particularly in compilation rate and bug detection. Based on the model information in Table 2, we find that newer models tend to generate higher-quality mutations, which aligns with the trends observed in other code-related tasks (Yang et al., 2024).
Finding 7 (RQ3): Two GPT-4 models perform best, while DeepSeek has a similar performance to GPT-3.5, standing out among open-source models. The newer models exhibit better performance in mutation generation.
4.4. RQ4: Impacts of Different Prompts
Table 7. Comparison Results of Different Prompts
Metric Category Metric GPT-3.5 DC-236b
P1 P2 P3 P4 P1 P2 P3 P4
—— Mut. Score 0.708 0.706 0.736 0.782 0.712 0.723 0.719 0.737
Mut. Count 48670 49126 42775 40844 53000 53177 52687 53130
Cost Avg. Gen. Time 1.79 1.78 1.80 1.87 4.25 4.25 4.20 4.40
Usability Compilability Rate 65.7% 63.8% 61.5% 60.6% 74.2% 75.1% 70.2% 72.6%
Useless Mut. Ratio 13.6% 11.3% 21.8% 14.8% 8.9% 3.8% 11.7% 9.8%
Eq. Mut. Rate 2.2% 1.9% 2.4% 2.5% 1.2% 1.3% 1.0% 1.5%
Behavior Real Bug Detectability 81.4% 80.7% 75.5% 68.0% 85.0% 83.7% 82.0% 82.7%
Coupling Rate 15.4% 15.1% 11.6% 11.7% 14.8% 14.6% 13.9% 14.4%
Ochiai Coefficient 66.3% 65.4% 63.1% 64.7% 66.7% 64.1% 64.4% 65.0%
@@ 4.4. RQ4: Impacts of Different Prompts
Table 7 presents the comparative results of GPT-3.5 and DeepSeek using different prompts listed in Section 3.5.3. Prompts P1 to P3 simplify progressively, each containing less information than its predecessor, while P4 is the most complex, enhancing P1 with test suite code.
@@
Overall, P1, the default prompt, excels in all Behavior Metrics on both models, indicating it can generate mutations more likely to resemble real bugs and effectively detect them. In addition, for all Usability Metrics, P1 achieves the second-best performance. P2, created by removing few-shot examples from P1, performs better in usability. P3, provided only with the code element to be mutated, achieves the lowest equivalent mutation rate. Conversely, P4, which extends P1 with additional test suite code, performs the worst in most cases.
@@
Finding 8 (RQ4): The prompt with the whole method and few-shot examples as context (i.e., P1) achieves the best performance across all Behavior Metrics, whereas adding the code of test suites (i.e., P4) decreases performance.
@@ 4.5. RQ5: Root Causes and Error Types of Non-Compilable Mutations
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 4.1, LLMs generate a significant number of non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT-3.5, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified 9 distinct error types, as shown in Table 8.
@@
Shown as Table 8, the most common error type, Usage of Unknown Methods, accounts for 27.3% of the total errors, revealing the hallucination issue of generative models (Huang et al., 2023). Code Structural Destruction is the second most common error, accounting for 22.9%, indicating that ensuring generated code is syntactically correct remains a challenge for current LLMs. This result suggests that there is still significant room for improvement in current LLMs.
Finding 9 (RQ5): The mutations generated by GPT models have nine main types of compilation errors, with Usage of Unknown Methods and Code Structural Destruction being the two most prevalent types.
@@
(b) GPT-4o
(c) GPT-4o-M
(d) DS-236b
(e) SC-16b
(f) CL-13b
(g) LEAM
(h) $\mu$Bert
(i) Major
@@
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major), the code locations with MemberReference and MethodInvocation are the top-2 prevent AST node types. For example, for GPT LLMs and DeepSeek, there about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 1.11% (DeepSeek), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% ($\mu$Bert), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation. Thus for LLMs, deletion is not the major reason for non-compilation.
Finding 10 (RQ5): Member references and method invocations within the code context are the most likely triggers for LLMs to generate non-compilable mutations.
@@ 5.1.1. Experiment Setup of the Context Length
QB-6a (Default) v.s. QB-3 0.9701 6.55E-05 0.9996 1.70E-10
QB-6a (Default) v.s. QB-3b 0.9762 3.31E-05 0.9984 1.00E-08
QB-6a (Default) v.s QB-9 0.9762 3.31E-05 0.9985 7.98E-09
QB-6a (Default) v.s D4J-6 0.9524 2.60E-04 0.9980 2.13E-08
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 9. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the upper part of Table 10. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the $p$-values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
@@ 5.1.2. Impact of Different Few-Shot Examples
We further compute their similarity, shown as the lower part of Table 10. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the $p$-values are all below 0.05. Therefore, we can conclude the performance of different prompts is very similar.